### v1

# ORBIT: On-policy Exploration-Exploitation for Controllable Multi-Budget Reasoning

Abstract
Recent Large Reasoning Models (LRMs) achieve strong performance by leveraging long-form Chain-of-Thought (CoT) reasoning, but uniformly applying overlong reasoning at inference time incurs substantial and often unnecessary computational cost. To address this, prior work explores various strategies to infer an appropriate reasoning budget from the input. However, such approaches are unreliable in the worst case, as estimating the minimal required reasoning effort is fundamentally difficult, and they implicitly fix the trade-off between reasoning cost and accuracy during training, limiting flexibility under varying deployment scenarios. Motivated by these limitations, we propose ORBIT, a controllable multi-budget reasoning framework with well-separated reasoning modes triggered by input. ORBIT employs multi-stage reinforcement learning to discover Pareto-optimal reasoning behaviors at each effort, followed by on-policy distillation to fuse these behaviors into a single unified model. Experiments show that ORBIT achieves (1) controllable reasoning behavior over multiple modes, (2) competitive reasoning density within each mode, and (3) integration of these frontier policies into a single unified student model while preserving clear mode separation and high per-mode performance.
ORBIT: On-policy Exploration-Exploitation
for Controllable Multi-Budget Reasoning
Kun Liang1,2,∗   Clive Bai3   Xin Xu3,4   Chenming Tang1,2,∗ Sanwoo Lee1,2,∗Weijie Liu3Saiyong Yang3,†Yunfang Wu1,2,† 1School of Computer Science, Peking University 2National Key Laboratory for Multimedia Information Processing, Peking University 3LLM Department, Tencent 4The Hong Kong University of Science and Technology ✉ kliang25@stu.pku.edu.cn   ✉ wuyf@pku.edu.cn
footnotetext: Work done during internship at Tencent.footnotetext: Corresponding Authors.
1 Introduction
Figure 1: ORBIT demonstrates controllable reasoning behavior comparable to advanced systems, covering multiple reasoning modes: low/mid/high for o3-mini and gpt-oss-120B, and an extended set of low/mid/high/extra-high for ORBIT.
Recent advances in Reinforcement Learning with Verifiable Rewards (RLVR) (Guo et al., 2025; Team et al., 2025; Xiaomi et al., 2025) have significantly pushed the frontier of complex task solving and long-form Chain-of-Thought (CoT) (Wei et al., 2022) reasoning in large language models. By encouraging explicit reasoning trajectories, these methods enable models to explore richer solution paths and achieve strong performance on challenging benchmarks (Guo et al., 2025). However, the continued push toward greater reasoning depth introduces a practical trade-off between reasoning effort and operational efficiency: increasingly long reasoning traces incur higher latency and computational cost, and are often unnecessary for simpler queries (Yang et al., 2025b).
Notably, this observation concerns not the use of long-chain reasoning during training, but its uncalibrated application at inference time, as longer reasoning chains often facilitate exploration of richer solution paths and contribute to higher peak accuracy (Setlur et al., 2025). The challenge primarily emerges at deployment time, where models trained to reason deeply may apply the same heavy reasoning patterns indiscriminately across inputs.
To optimize inference-time efficiency, prior work has explored various strategies for adjusting reasoning effort according to task demands. (Luo et al., 2025a; Lou et al., 2025; Luo et al., 2025b; Zhang et al., 2025). Across these approaches, the underlying goal is effectively the same: to estimate an optimal allocation of reasoning effort.
However, this estimation problem is fundamentally unreliable. Determining the minimal amount of reasoning required to solve an instance is closely related to estimating its Kolmogorov complexity (Li and Vitnyi, 2008), which is known to be uncomputable and connected to the halting problem theoretically (Graves, 2017). Empirically, adaptive mechanisms can also misallocate reasoning effort—truncating reasoning on instances that would benefit from deeper deliberation (Aggarwal et al., 2025; Jin et al., 2025), while preserving unnecessarily long reasoning for simpler instances (Chen et al., 2025a).
Moreover, these inference-agnostic approaches implicitly determine the trade-off between reasoning effort and expected accuracy during training, reflecting a fixed preference over cost and performance. While suitable for standardized benchmarks, this design may be less flexible in deployment scenarios where user preferences and system constraints vary across queries.
Motivated by these limitations, we rethink a controllable reasoning model that can flexibly adjust its reasoning effort according to user preferences at inference time, meanwhile achieving competitive reasoning densities within each effort.
We introduce ORBIT, a fully on-policy training framework for controllable multi-budget reasoning. By exploring under different training context windows, ORBIT identifies a Pareto frontier of reasoning behaviors and effectively integrates these into a single model whose discrete reasoning modes can be triggered by prompts. As illustrated in Figure 1, ORBIT achieves clear and well-separated reasoning modes similar to state-of-the-art controllable systems, while maintaining strong per-mode accuracy and token-efficient reasoning.
Our contributions are as follows:
  1. 1.
    We propose ORBIT, a fully on-policy training framework for controllable multi-budget reasoning, enabling a single model to generate responses across multiple reasoning modes triggered by mode-specific queries.
  2. 2.
    We introduce an expansion-compression exploration strategy that efficiently discovers mode-specific reasoning frontiers under varying context constraints, yielding high-accuracy policies at each reasoning effort.
  3. 3.
    We design a multi-teacher on-policy distillation pipeline with mode-aware initialization, which fuses these frontier policies into a single unified student model while preserving strict mode separation and high performance across all modes.
2 Preliminary
Problem Statement
. We study controllable multi-budget reasoning, aiming to learn a single, unified policy $\pi_{\phi}(o\mid q,c)$ that generates distinct, mode-specific reasoning behaviors for a query $q\in\mathcal{D}$, where the mode $c\in\mathcal{C}=\{c_{1},\dots,c_{K}\}$ can be selected or triggered at inference time, corresponding to different reasoning effort.
Crucially, each induced behavior is expected to be strong in its own right: for every mode $c_{k}$, the corresponding policy $\pi_{\phi}(\cdot\mid q,c_{k})$ should achieve high task accuracy relative to its reasoning budget, rather than representing a compromised linear interpolation performance.
The core challenge, therefore, lies in learning such a unified yet controllable policy under a fixed training budget, while maintaining clear behavioral separation across modes.
RLVR
Following Ouyang et al. (2022) and Shao et al. (2024), we model language generation as autoregressive sampling from a conditional probability distribution $\pi_{\theta}(y|x)$, and our training goal is to maximize the expected correctness reward $r(q,o)\in\{0,1\}$:
$\max_{\theta}\mathbb{E}_{q,o\sim\pi_{\theta}}[r(q,o)],$ (1)
where $q$ represents the input query and $o$ represents the output response. We adopt RLVR as our training method for two reasons: (i) RLVR provides an effective way to improve reasoning capability by directly reinforcing self-generated sequences with correctness feedback; (ii) In autoregressive models, modifying intermediate tokens in an offline manner—especially in long-form CoT reasoning—can be prone to significant distribution shifts (Xie et al., 2025), whereas on-policy RLVR preserves distributional consistency by training on complete self-generated sequences.
In this work, we adopt Group Relative Policy Optimization (GRPO) (Shao et al., 2024) as our RLVR algorithm for simplicity. Unlike traditional methods (Schulman et al., 2017), GRPO bypasses the value model by computing advantages relative to a group of $G$ responses $\{o_{1},\dots,o_{G}\}$ sampled from $\pi_{\theta_{\text{old}}}$. The objective is:
$\mathcal{J}_{\text{GRPO}}(\theta)=\mathbb{E}_{q\sim\mathcal{D}}\Biggl[\frac{1}{G}\sum_{i=1}^{G}\frac{1}{|o_{i}|}\sum_{t=1}^{|o_{i}|}\\ \min\Bigl(r_{i,t}(\theta)\hat{A}_{i,t},\text{clip}(r_{i,t}(\theta),1-\epsilon,1+\epsilon)\hat{A}_{i,t}\Bigr)\Biggr],$ (2)
where the importance ratio $r_{i,t}(\theta)$ and the group-relative advantage $\hat{A}_{i}$ are defined as:
$r_{i,t}(\theta)=\frac{\pi_{\theta}(o_{i,t}|q,o_{i,<t})}{\pi_{\theta_{\text{old}}}(o_{i,t}|q,o_{i,<t})},$ (3)
$\hat{A}_{i}=\frac{r(q,o_{i})-\text{mean}(\{r(q,o_{j})\})}{\text{std}(\{r(q,o_{j})\})}.$ (4)
3 Methodology
Figure 2: The ORBIT framework overview: (1) In the Exploration stage, multi-stage RL discovers specialized reasoning frontiers under varying context from the expansion-compression loop. (2) In the Exploitation stage, these behaviors are unified into a single student model via model merging and multi-teacher on-policy distillation.
We approach controllable multi-budget reasoning through an exploration–exploitation perspective. The challenge is twofold: for each reasoning mode, the model must achieve high task performance within a specified reasoning budget, while simultaneously supporting multiple modes within a single model.
To address this, we decompose the problem into two complementary phases. In the Exploration phase, multi-stage RL places computational effort in generating sequences in semantic space for exploring under varying context budgets, thus discovering high-quality per-mode strategies. In the Exploitation phase, we consolidate these discovered policies into a single student model via on-policy distillation, preserving mode-specific behaviors while enabling a single efficient model for all reasoning modes. Figure 2 provides an overview of this framework.
3.1 Expansion-Compression Loop for Mode-Specific Policy Exploration
During exploration, each reasoning mode $c\in\mathcal{C}$ is subject to a context budget reflecting its intended reasoning effort. The training objective is to discover high-quality strategies that maximize task accuracy under this budget. To achieve this, we adopt the simple length regularizing method–truncation along with a scheduling strategy and formalize the exploration stage as traversal along the accuracy–length Pareto frontier.
Dual-Objectivity under Truncation
For a given context budget $L_{k}$ and reasoning mode $c\in\mathcal{C}$, we consider a policy $\pi_{\theta_{k}}$ and define its truncated, mode-specific version:
$\tilde{\pi}_{\theta_{k}}^{(c)}(o;L_{k}):=\pi_{\theta_{k}}(o\mid q,c)\,\mathbb{I}(|o|\leq L_{k}),$ (5)
For a given mode $c$, the expected reward (accuracy) under truncation is
$J^{(c)}(\pi_{\theta_{k}})=\mathbb{E}_{q\sim\mathcal{D}}\mathbb{E}_{o\sim\tilde{\pi}_{\theta_{k}}^{(c)}(\cdot\mid q)}[r(q,o)],$ (6)
We define the non-truncation probability
$p_{\text{non-trunc}}^{\pi,(c)}(q):=\mathbb{P}_{o\sim\pi_{\theta_{k}}(\cdot\mid q,c)}(|o|\leq L_{k}),$ (7)
and the corresponding renormalized truncated policy
$\pi_{\theta_{k}}^{T_{k},(c)}(o\mid q):=\frac{\pi_{\theta_{k}}(o\mid q,c)\,\mathbb{I}(|o|\leq L_{k})}{p_{\text{non-trunc}}^{\pi,(c)}(q)}.$ (8)
Then the truncated expected reward decomposes as
$J^{(c)}(\pi_{\theta_{k}})$ $=\mathbb{E}_{q\sim\mathcal{D}}\Big[p_{\text{non-trunc}}^{\pi,(c)}(q)\,\bar{r}_{\text{non-trunc}}^{\pi,(c)}(q)\Big],$ (9)
$\bar{r}_{\text{non-trunc}}^{\pi,(c)}(q)$ $:=\mathbb{E}_{o\sim\pi_{\theta_{k}}^{T_{k},(c)}(\cdot\mid q)}[r(q,o)].$ (10)
This decomposition shows that, for each mode $c$, optimization under truncation involves two coupled directions:
  • increasing $p_{\text{non-trunc}}^{\pi,(c)}$, i.e., concentrating probability mass on valid-length trajectories;
  • increasing $\bar{r}_{\text{non-trunc}}^{\pi,(c)}$, i.e., improving the semantic quality of non-truncated trajectories.
Harvesting Along the Frontier via an Expansion–Compression Loop
To efficiently obtain high-quality per-mode strategies under truncation, we introduce a one-time Expansion–Compression Loop for scheduling context window in RLVR training:
  • Expansion (Semantic Optimality). We first enlarge the context budget to $L_{\mathrm{exp}}\gg\mathbb{E}[|o^{+}|]$, where $o^{+}$ denotes correct, within-window trajectories, ensuring truncation is inactive and the policy focuses on semantic correctness. This allows exploration of harder instances and establishes a global performance ceiling.
  • Sequential Compression (Density Optimality).
    Starting from the expanded policy, we progressively tighten the context budget $L_{k}$. This constraint imposes truncation, inducing a discontinuous length distribution that can destabilize training through two mechanisms: (i) increased reward variance due to the scarcity of valid-length correct trajectories, which leads to noisy gradient estimates; and (ii) “survival mode,” where the model collapses to the shortest, trivial solutions, suppressing the complex reasoning required for harder instances. To balance the need for mode separability with training stability, we adopt a logarithmic compression schedule (i.e., $L_{k+1}=L_{k}/2$). Furthermore, we train each stage to saturation—monitored via a Monte Carlo estimate of the truncation ratio. This ensures that the policy remains locally stable, maintains semantic quality, and preserves a sufficient density of valid trajectories before the next compression.
The loop yields stage-wise, Pareto-optimal per-mode experts along the accuracy–length frontier, which then serve as teachers for on-policy distillation to consolidate all modes into a single model.
3.2 On-policy Distillation for Frontier Policy Exploitation
While the exploration stage identifies frontier policies, REINFORCE-style updates provide only a sparse $O(1)$ supervision signal per rollout sequence (uniform $\hat{A}$ across all tokens). Such scalar feedback is insufficient not only for efficient consolidation, but also for fine-grained, token-level reasoning behaviors anchoring across multiple mode-specific experts.
[…]
On-Policy Distillation Mode Fusing
Common intuition trains distillation with an offline dataset with query and responses sampled from the teacher model and train the student in an SFT manner.
However, distilling auto-regressive student models in this way leads to a training–inference distribution mismatch, commonly known as exposure bias, because supervision is provided on teacher-generated trajectories that the student policy may never enter during inference, preventing it from learning to correct its own self-induced errors (Agarwal et al., 2023).
Based on these intuition, we adopt On-Policy Distillation (Agarwal et al., 2023; Lu and Lab, 2025), minimizing the Reverse Kullback-Leibler (RKL) divergence:
$\mathcal{L}_{\text{distill}}(\phi)=\mathbb{E}_{x\sim\mathcal{D},y\sim\pi_{\phi}(\cdot|x,p_{k})}\left[\log\frac{\pi_{\phi}(y|x,p_{k})}{\theta^{(k)}(y|x,p_{k})}\right].$ (11)
RKL is inherently mode-seeking, encouraging the student $\pi_{\phi}$ to focus on the high-density reasoning modes of the expert $\theta^{(k)}$ dictated by prompt $p_{k}$. By concentrating probability mass on these modes, the student is less likely to assign weight to irrelevant trajectories or non-target modes. This ensures that the reasoning knowledge from each expert is effectively consolidated, while preserving the model’s broader generative capabilities.
Formally, the objective function is defined as:
$\mathcal{J}_{\text{Distill}}(\phi)=\mathbb{E}_{k\sim U(1,K)}\mathbb{E}_{\begin{subarray}{c}q\sim\mathcal{D}\\ o\sim\pi_{\phi}(\cdot|q,p_{k})\end{subarray}}\\ \left[\sum_{t}\log\frac{\pi_{\phi}(o_{t}|q,p_{k},o_{<t})}{\pi_{\theta^{(k)}}(o_{t}|q,o_{<t})}\right],$ (12)
Mode-aware initialization via Model Merging
Despite the theoretical advantages of RKL, on-policy distillation suffers a "cold-start": if the student $\pi_{\phi}$ initially lacks confidence in the target modes, rollouts rarely reach the expert’s high-density manifold, limiting the supervisory signal. We mitigate this by initializing the student via model merging, combining the stage-wise frontier models $\{\pi_{\theta^{(k)}}\}_{k=1}^{K}$ obtained during exploration, for aligning input-output mappings with the teacher $\pi_{\theta^{(k)}}$. This provides a more stable start and accelerating convergence under our RKL distillation setting; details can be found in Appendix B.
4 Experiments
4.1 Experimental Settings
We implemented ORBIT on three models— DeepSeek-Distill-Qwen-1.5B, Qwen3-4B-2507-Thinking, and Openmath-Nemotron-7B—each configured with four reasoning modes: Low, Mid, High, and Xhigh. The training context lengths for these modes were set according to the model: 2K, 4K, 8K, and 16K tokens for DeepSeek-Distill-Qwen-1.5B, and 4K, 8K, 16K, and 32K tokens for both Qwen3-4B-2507-Thinking and Openmath-Nemotron-7B. Specifically, DeepSeek-Distill-Qwen-1.5B was trained on the DAPO-Math-17K dataset (Yu et al., 2025), while Qwen3-4B-2507-Thinking and Openmath-Nemotron-7B were trained on Polaris-53K (An et al., 2025). Notably, the training procedure is scheduled in reverse order. Further implementation and training details are provided in Appendix C.
Baselines. We first evaluate our full pipeline against the recent controllable reasoning models, including o3-mini (OpenAI, 2025b), gpt-oss-120b (OpenAI, 2025a), and ThinkDial (He et al., 2025). Next, we compare ORBIT’s within-mode performance against current RL-based adaptive and efficient reasoning methods (Aggarwal and Welleck, 2025; Arora and Zanette, 2025; Zhang et al., 2025; Fang et al., 2025; Tu et al., 2025; Liu et al., 2025) on DeepSeek-Distill-Qwen-1.5B. Finally, we ablate ORBIT with other potentially used methods and baselines.
Evaluation Benchmarks. We evaluate our methods on three mathematical reasoning tasks: AIME 2024 & 2025, BeyondAIME (ByteDance-Seed, 2025), as well as representative Multi-task benchmarks: GPQA-Diamond (Rein et al., 2024), and MMLU-Pro (Wang et al., 2024).
Mathematics Multi-Task Overall
Models AIME 24 AIME 25 Beyond AIME GPQA MMLU
avg@32 Toks avg@32 Toks avg@8 Toks avg@8 Toks avg@1 Toks Avg. Toks
DeepSeek-Distill-Qwen-1.5B
Initial Model 28.6 18.8 22.8 18.5 10.1 18.6 28.0 12.7 35.5 7.2 25.0 15.2
- ORBIT Low 28.7 2.7 21.1 2.3 9.9 1.9 34.8 1.8 38.1 1.4 26.5 2.0
- ORBIT Mid 34.5 5.6 26.0 4.9 12.6 3.6 34.3 2.6 38.2 2.0 29.1 3.7
- ORBIT High 38.8 7.7 30.2 7.4 15.6 7.1 41.0 4.9 41.7 3.4 33.5 6.1
- ORBIT Xhigh 42.6 8.5 31.1 7.8 15.9 7.7 41.8 6.4 41.2 4.4 34.5 7.0
Qwen3-4B-2507-Thinking
Initial Model 81.6 20.1 77.1 18.3 52.4 27.6 66.1 8.8 74.4 4.4 70.3 15.8
- ORBIT Low 59.7 3.3 46.8 3.8 31.4 3.4 60.5 2.0 71.7 1.1 54.0 2.7
- ORBIT Mid 69.5 6.6 61.0 7.5 40.6 7.4 62.9 3.1 72.8 1.9 61.4 5.3
- ORBIT High 75.0 9.6 69.2 11.0 48.3 11.8 65.2 4.6 73.5 2.8 66.2 8.0
- ORBIT Xhigh 82.6 15.5 78.2 17.4 53.5 20.7 69.5 8.4 74.5 5.4 71.7 13.5
Openmath-Nemotron-7B
Initial Model 71.6 11.9 59.9 13.7 41.8 13.6 32.5 8.5 42.4 5.7 49.6 10.7
- ORBIT Low 60.0 5.2 36.5 5.7 31.5 4.7 31.5 2.4 39.3 1.8 39.8 4.0
- ORBIT Mid 66.4 7.4 45.2 8.3 35.0 7.5 32.3 4.2 41.5 2.9 44.1 6.1
- ORBIT High 71.4 10.0 57.1 11.6 41.4 10.9 31.9 6.1 41.7 4.4 48.7 8.6
- ORBIT Xhigh 72.0 11.0 57.8 12.1 41.8 12.1 34.8 6.6 42.2 5.0 49.7 9.4
Table 1: Comparison of reasoning modes of ORBIT across various reasoning tasks. “Avg@k” denotes the average accuracy (in %) over k generations (i.e., pass@1), and "Toks" indicates the average output length in thousands of tokens (K). Darker colors in the cell background denote better results.
4.2 Main Results
ORBIT Achieves Controllable Reasoning Behavior Across Domains
Table 1 illustrates the reasoning behaviors learned by ORBIT under different reasoning budgets across domains. We observe a clear and consistent separation between reasoning modes, with each exhibiting a stable reasoning density and a monotonic trade-off between reasoning effort and task performance. This separation indicates that ORBIT successfully learns query triggered inference-time separable distinct reasoning policies.
Importantly, these controllable behaviors transfer reliably across domains. When transferring from mathematical reasoning to general question answering benchmarks, the ordinal ordering of reasoning modes is preserved, while only minimal stepwise performance degradation is incurred.
ORBIT Achieves Competitive Reasoning Density Within Mode
Models AIME 24 AIME 25 Beyond AIME GPQA MMLU Overall
avg@32 Toks avg@32 Toks avg@8 Toks avg@8 Toks avg@1 Toks Avg. Toks
Initial Model 28.6 18.8 22.8 18.5 10.1 18.6 28.0 12.7 35.5 7.2 25.0 15.2
1) TLMRE-DS-1.5B ($\alpha=0.1$) 28.1 14.7 22.1 14.2 8.4 12.7 30.8 7.8 34.3 4.7 24.7 10.8
2) AdaptThink-1.5B ($\delta=0.05$) 25.2 7.6 21.3 8.1 9.9 5.1 31.3 6.7 36.9 4.1 24.9 6.3
3) AutoThink-DS-1.5B-Stage3 31.0 11.0 22.8 10.6 11.0 9.4 31.4 7.6 36.3 4.5 26.5 8.6
4) Laser-DE-L4096-1.5B 30.6 8.2 22.0 7.5 10.3 7.2 34.6 3.4 37.8 2.3 27.1 5.7
5) AutoThink-Stage3 39.5 9.0 29.0 9.0 12.4 8.0 35.7 4.7 39.6 3.6 31.2 6.9
6) L1-1.5B-Max 23.6 3.4 22.3 3.1 10.4 3.1 33.0 1.2 39.5 1.4 25.8 2.4
7) Thinkless-1.5B-RL 27.8 11.1 21.8 11.1 11.1 11.6 28.9 12.8 35.2 6.6 25.0 10.6
- ORBIT Low 28.7 2.7 21.1 2.3 9.9 1.9 34.8 1.8 38.1 1.4 26.5 2.0
- ORBIT Mid 34.5 5.6 26.0 4.9 12.6 3.6 34.3 2.6 38.2 2.0 29.1 3.7
- ORBIT High 38.8 7.7 30.2 7.4 15.6 7.1 41.0 4.9 41.7 3.4 33.5 6.1
- ORBIT Xhigh 42.6 8.5 31.1 7.8 15.9 7.7 41.8 6.4 41.2 4.4 34.5 7.0
Table 2: Comparison of reasoning modes of ORBIT trained from DeepSeek-Distill-Qwen-1.5B with efficient reasoning baselines across various reasoning tasks. Darker colors in the cell background denote better results.
We evaluate ORBIT on DeepSeek-Distill-Qwen-1.5B, comparing its within-mode accuracy–length trade-off against RL-based efficient reasoning methods.
As shown in Table 2, ORBIT generally achieves competitive trade-offs, often attaining higher accuracy for a fixed reasoning budget or requiring fewer reasoning tokens for a fixed accuracy level. Notably, these results are obtained with a simple truncation-based reward combined with our scheduling and fusion strategy, highlighting the effectiveness and simplicity of ORBIT.
5 Analysis
We conducted a series of analytical experiments on the best-performing Qwen3-4B-2507-Thinking.
Figure 3: Step-wise performance of ORBIT and joint RL on AIME24, with steps aligned by total training tokens for fair comparison. ORBIT starts from step 278, where OPD training begins. Sequential compression stages are omitted for clarity.
Comparison with END-to-END Joint RL Training
We first investigate whether mode separation can be integrated into the RL training, where we jointly train each single prompt under four different reasoning modes, each associated with a specific system prompt and corresponding context window. To ensure the peak performance, we start our joint RL training after expansion stage, and stabilize the training with rejection sampling to prevent training oscillation under the most aggressive truncation (e.g. 32K $\rightarrow$ 4K). All other training settings are identical to those of our sequential counterpart.
As shown in Figure 3, joint RL exhibits a concentrated behavior, with Mid,High,and Xhigh modes to convergence on same point.
We identify two key factors contribute to this. First, supervision from truncated sequences dominates optimization, consistent with our truncation-rate decomposition in Sec 3. Second, High mode sequences can satisfy their truncation requirements by compressing below low-mode levels. Both reasons cause a reward hacking on response length against a mode selective reasoning behavior.
Analysis of Fusing Strategies
We examine ORBIT’s fusing strategy by replacing on-policy fusion with offline distillation, where consolidation proceeds without additional on-policy rollouts, relying instead on a pre-sampled dataset from the exploration-stage frontier models. Both variants are initialized from the same merged model to minimize behavioral inconsistency.
As shown in Figure 4, on-policy fusion achieves slightly higher performance than offline distillation, while overall stability and convergence remain comparable. This outcome can be attributed to the strong framework-level alignment in ORBIT: all teacher and student models share a sequential training with same base model trained on same data with same algorithm, effectively reducing policy mismatch and rendering offline distillation in this setting closer to a semi-offline procedure.
Beyond performance, on-policy fusion provides operational advantages: it generates rollouts dynamically, adapts convergence on-the-fly, and integrates naturally with RL-based correctness signals. In contrast, offline distillation decouples dataset construction from training, allowing for scalable evaluation via pre-collected datasets, which can be leveraged in techniques like Rejection Fine-tuning. The choice between these strategies, or their combination, can be tailored depending on specific application requirements.
Figure 4: Comparison of OPD and offline distillation as the fusing strategy on AIME24, with steps aligned by total training samples for fair comparison.On-policy fusion achieves slightly higher performance while both strategies show similar trend.
ORBIT’s Soft-Mode Constraint Elicits Cognition Separation
Figure 5: Cognition analysis of ORBIT across reasoning modes, vertical axis denotes average count per sample.
To further analyze how ORBIT’s mode control shapes reasoning behavior, we evaluate it on MATH500 (Lightman et al., 2023) and visualize the evolution of reasoning components in Figure 5. We depict ORBIT’s reasoning behavior along three components: Information Extraction, Logical Deduction, and Self-Reflection—each capturing a distinct aspect of the model’s cognitive process. As reasoning modes increase, Logical Deduction steps grow roughly linearly, reflecting the model’s structured problem-solving growth, calibrated by training context and demonstrating clear predictability across modes. More importantly, reasoning control elicits an apparent System-2 thinking pattern via an exponential-like increase of self-reflection, increasing computational effort toward high-level reflective reasoning as modes advance. Meanwhile, Information Extraction remains nearly constant, showing the model avoids inflating token usage through redundant condition repetition. Together, these trends show ORBIT elicits hierarchical cognitive behavior, where different reasoning components occupy distinct functional levels and higher reasoning modes systematically amplify contributions of more reflective, high-level processes.
6 Related Work
Reasoning Effort Adjustment in LRMs
Existing works on reasoning effort adjustment primarily target overthinking through RL-based length penalties (Aggarwal and Welleck, 2025; Arora and Zanette, 2025; Yi et al., 2025; Zhang et al., 2025; Fang et al., 2025; Lou et al., 2025), verification training (Chen et al., 2025b), early exit strategies (Yang et al., 2025a; Jiang et al., 2025), or inference-time interventions (Wang et al., 2025). Few lines of methods tackle the underthinking problem by decoding time intervention (Muennighoff et al., 2025; Jin et al., 2025). ORBIT try to tackle these problems by offering flexible reasoning modes at inference time.
Controllable Reasoning Generation
Prior work for reasoning control can be boradly categroized in two main streams (1) Hard Token Budget Constraint: A common approach enforces reasoning efficiency through externally specified constraints, most notably by fixing explicit token budgets that cap the length of intermediate reasoning Muennighoff et al. (2025); Sun et al. (2025); Aggarwal and Welleck (2025); Anthropic (2025), relying on an absolute tokenwise budget estimation across all inputs. (2) Discrete-Mode Systems: offer an intuitional discrete reasoning levels, enabling users to explicitly specify their preference for the efficiency-performance trade-off OpenAI (2025a); He et al. (2025), representing a soft form of control based on model’s own capability.
7 Conclusion
In this paper, we propose ORBIT, a fully on-policy training framework for controllable reasoning models that operates at a target reasoning level while maintaining high reasoning density under a given context. We leverage a simple yet effective truncation reward and adopt a one-time expansion-compression loop strategy to explore the reasoning frontier. On-policy distillation is then applied to consolidate multiple modes into a single unified model, initialized from merged frontier policies and enabling mode selection via natural language prompts during training. Experiments across LRMs of different sizes demonstrate that ORBIT achieves controllable reasoning behavior and delivers competitive accuracy–length performance compared to existing RL-based efficient reasoning methods within each mode.
Limitations
We elucidate the limitations of this work as follows: First, in sequential RLVR exploration training, the binary context partitioning used in compression stage may not be the best scheduling strategy for balancing mode separation in test-time and training stabilization. On the other hand, for a highly fixed model with output entropy below a certain threshold, on-policy RL-based training may suffer from low sample efficiency, and in extreme cases, the training signal may be too weak to achieve desirable saturation, limiting compression progress. Second, long-form RL-based training can be computationally expensive for larger models, as each rollout requires auto-regressive sampling, which may limit scalability to very large architectures or long-context reasoning tasks.
Ethical Considerations
Use of AI Assistants
The algorithmic design and main methodology presented in this work were developed through human-led research and reasoning. During implementation, we utilized GitHub Copilot 3111https://github.com/features/copilot for coding assistance. We affirm that the primary content and logic of the code are entirely our own work.
Potential Risks
This work introduces ORBIT (On-policy Exploration-Exploitation for Controllable Multi-Budget Reasoning), a framework for training language models with controllable reasoning modes. We acknowledge potential risks associated with controllable reasoning models. Misuse could lead to over-reliance on automated reasoning, propagation of incorrect conclusions, or reinforcement of biases present in the training data. Moreover, the ability to choose reasoning modes in ORBIT may lead to misleading outputs if used improperly or adversarially.
Appendix A System Prompts For Each Reasoning Mode
To build a semantically meaningful guidance for discrete mode separated reasoning behavior, we follow Yang et al. (2025b) and generalized to four mode, where each one of them represent semactically aligned expected reasoning behavior under corresponding mode.
/* Low Mode */ You have extremely limited time to think and respond to the user’s query. Every additional second of processing and reasoning incurs a significant resource cost, which could affect efficiency and effectiveness. Your task is to prioritize speed without sacrificing essential clarity or accuracy. Provide the most direct and concise answer possible. Avoid unnecessary steps, reflections, verification, or refinements UNLESS ABSOLUTELY NECESSARY. Your primary goal is to deliver a quick, clear and correct response. /* Medium Mode */ You have sufficient time to think and respond to the user’s query, allowing for a more thoughtful and in-depth answer. However, be aware that the longer you take to reason and process, the greater the associated resource costs and potential consequences. While you should not rush, aim to balance the depth of your reasoning with efficiency. Prioritize providing a well-thought-out response, but do not overextend your thinking if the answer can be provided with a reasonable level of analysis. Use your reasoning time wisely, focusing on what is essential for delivering an accurate response without unnecessary delays and overthinking. /* High Mode */ You have a substantial amount of time to think and respond. The pressure on reasoning cost is low, allowing you to prioritize accuracy and robustness over speed. You should verify your key steps and consider alternative interpretations to ensure your answer is reliable. While you are encouraged to think deeply, aim for a confident solution rather than an endless exploration. If you encounter ambiguity, take the time to resolve it, but remain focused on deriving the correct conclusion efficiently within a generous budget. /* Xhigh Mode */ You have unlimited time to think and respond. There are absolutely no constraints on reasoning length or associated costs. Your singular goal is to achieve the highest possible level of accuracy and certainty. You must exhaustively explore the problem space, considering all edge cases, theoretical nuances, and alternative methodologies. Engage in rigorous, iterative self-correction and deep verification to eliminate even the slightest possibility of error. Do not settle for a ’likely’ correct answer; explore every angle until you are certain. Perfection is the only standard.
Table 3: System prompts used for triggering Low, Medium, High, Xhigh modes
Mathematics Multi-Task Overall
Models AIME 24 AIME 25 Beyond AIME GPQA MMLU
avg@32 Toks avg@32 Toks avg@8 Toks avg@8 Toks avg@1 Toks Avg. Toks
DeepSeek-Distill-Qwen-1.5B
Initial Model 28.6 18.8 22.8 18.5 10.1 18.6 28.0 12.7 35.5 7.2 25.0 15.2
- Teacher Low 28.3 2.6 19.0 2.0 11.0 2.0 34.3 1.8 37.6 1.4 26.0 2.0
- Teacher Mid 35.2 4.9 26.4 4.5 10.4 3.6 32.9 2.9 38.2 2.2 28.6 3.6
- Teacher High 40.0 7.0 29.8 6.7 14.0 6.3 39.8 4.8 40.7 3.8 32.9 5.7
- Teacher Xhigh 42.5 10.7 32.9 11.0 16.0 11.7 39.5 8.3 40.0 7.7 34.2 9.9
Qwen3-4B-2507-Thinking
Initial Model 81.6 20.1 77.1 18.3 52.4 27.6 66.1 8.8 74.4 4.4 70.3 15.8
- Teacher Low 64.4 5.0 54.3 5.5 36.6 5.2 63.9 3.0 71.5 0.9 58.1 3.9
- Teacher Mid 66.3 5.9 57.7 6.4 40.0 6.4 62.9 32.9 72.3 1.9 59.8 10.7
- Teacher High 75.1 9.4 69.1 10.6 46.0 11.7 65.4 4.9 73.4 3.0 65.8 7.9
- Teacher Xhigh 83.2 16.8 78.2 19.0 52.7 22.7 67.1 7.9 74.1 5.8 71.1 14.4
Openmath-Nemotron-7B
Initial Model 71.6 11.9 59.9 13.7 41.8 13.6 32.5 8.5 42.4 5.7 49.6 10.7
- Teacher Low 55.2 4.5 33.1 4.5 29.8 3.7 29.3 2.2 39.2 1.7 37.3 3.3
- Teacher Mid 67.1 6.7 44.5 7.6 36.3 6.6 32.4 4.4 41.2 3.0 44.3 5.7
- Teacher High 74.3 11.1 62.1 12.9 41.9 12.1 30.9 7.9 42.6 5.3 50.4 9.9
- Teacher Xhigh 75.9 13.2 65.1 14.9 43.8 15.0 33.7 9.8 43.1 6.5 52.3 11.9
Table 4: Comparison of reasoning modes of Teacher across various reasoning tasks.
Appendix B Mode-aware initialization via Model Merging
As described in section3, optimizing Eq. 12 in practice faces a "cold-start" challenge. To mitigate this misalignment, we proposed using model merging as initialization of the student model.
Model merge act as a simple linear combination of model parameters. Numerous methods have been proposed, we use average merging (Wortsman et al., 2022) here for simplicity. This process can be defined as follows:
$\pi_{\phi}=\frac{1}{K}\sum_{k=1}^{K}\pi_{\theta^{(k)}}$ (13)
We compared the merged model and the final policy model of sequentially multi-stage RL stage as initialization, since only the two models have been trained under all reasoning modes. We first use Accmulated Misalignment Rate(AMR), which depicted whether the top-1 token probability is same between two models, greatly consistent with on-policy distillation process, for a intuitive behavioral visualization. Defined as:
$\text{AMR}(\pi,\theta^{(k)})=\mathbb{E}_{q\sim\mathcal{D},o\sim\pi(\cdot|q)}\left[\frac{1}{|o|}\right.\\ \left.\sum_{t=1}^{|o|}\mathbb{I}\left(\arg\max\pi(o_{t}|\cdot)\neq\arg\max\pi_{\theta^{(k)}}(o_{t}|\cdot)\right)\right]$ (14)
Figure 6: Misalignment rate on merged model generated and low mode model generated sequences
Figure 7: RKL training loss dynamic for different initialization in OPD
As shown in Figure 6, this effort-aware initialization significantly reduces the initial misalignment gap, further validated to providing a more stable starting point and accelerates the convergence of the subsequent RKL distillation.
We further examine the training performance under different initialization, as shown in Figure 7, merged model consistently outperforms final stage policy by a lower per token RKL penalty, demonstrating the effectiveness of such initialization.
Appendix C Training Details
We build on the VeRL codebase (Sheng et al., 2024) with RLVR. Following the DAPO recipe (Yu et al., 2025), which allows for higher clipping—a practice that has been widely adopted in RL training, we also use rejection sampling as a further ensurance for training stability.
Training Configuration. We maintain consistent hyperparameters throughout the exploration RL stage, using a batch size of $256$, rejection pools of $512$, and a learning rate of $1\times 10^{-6}$ without warm-up. For generation, we apply standard rollout settings: temperature $1$, top-p $1$, top-k 1-1, and $8$ rollouts per problem.
We train DeepSeek-Distill-Qwen-1.5B using the DAPO-Math-17K dataset (Yu et al., 2025), while Qwen3-4B-2507-Thinking and Nemotron-7B are trained on the Polaris-53K (An et al., 2025).
For exploitation, we train OPD with a batch size of $1024$, one generation per sample, and a token-wise KL penalty coefficient of $1.0$.
Context Window Scheduling. We implement an expansion-compression loop strategy to dynamically manage context windows in Exploration stage RL. Specifically, the maximum output length follows a schedule of $32\text{K}\rightarrow 16\text{K}\rightarrow 8\text{K}\rightarrow 4\text{K}$ for Qwen3-4B-2507-Thinking and Openmath-Nemotron-7B, and $16\text{K}\rightarrow 8\text{K}\rightarrow 4\text{K}\rightarrow 2\text{K}$ for DeepSeek-Distill-Qwen-1.5B. In both cases, the first stage of the context window serves as the expansion phase.
Appendix D More results On Frontier Model’s Performance
We provide the evaluation results for each teacher model from exploration stage in Table 4.